### v1

# Lollipops, dense cycles and chords

Abstract
We prove that every graph $G$ with minimum degree at least $k\geq 2$ contains a cycle $C$ that is dense in several respects. First, it has at least $k+1$ vertices each having at least $k$ neighbors in $C$ (so $C$ has at least $\frac{(k+1)(k-2)}{2}$ chords). The inequality is tight and provides a sharp upper bound for the chromatic number of graphs where all cycles have less than $\ell$ chords for all $\ell\geq 0$. Then, we show that some edges of $C$ can be contracted to obtain a graph with high minimum degree (we call such a minor of $C$ a cyclic minor). We then investigate further cycles having cliques as cyclic minors, and show that minimum degree at least $O(k^{2})$ guarantees a cyclic $K_{k}$-minor.
1 Introduction
Many theorems in graph theory state that a sufficiently high minimum degree in a graph guarantees the existence of some substructure that is in some sense dense, complex or well connected. We list below some classical examples:
  • a highly connected subgraph [Mad72];
  • a large clique as a minor [Kos82, dlV83];
  • a large clique as a topological minor [BT98];
  • a large biclique as a subgraph or a subdivision of some prescribed graph as an induced subgraph [KO04] and
  • a $k$-linked subgraph [TW05].
Here, we add some items to this list, by exhibiting cycles that are dense in several ways as we explain now.
Many Chords
Our first result is the following.
Theorem 1.
If $G$ has minimum degree at least $k\geq 2$, then $G$ contains a cycle $C$ containing at least $k+1$ vertices each having at least $k$ neighbors in $C$ (so $C$ has at least $\frac{(k+1)(k-2)}{2}$ chords). Moreover, by deleting all vertices not contained in $C$ and contracting some of the edges of $C$, we can obtain
  • a graph of minimum degree at least $\left\lceil\frac{k+2}{2}\right\rceil$, and
  • a graph of average degree at least $\tfrac{2}{3}(k+1)$.
It is worth noting that contrary to the theorems mentioned in the short survey above, some conclusions of 1 are tight, which can be easily seen by considering a complete graph $K_{t}$, that has minimum degree $k=t-1$, and that obviously contains a cycle with exactly $k+1$ vertices of degree exactly $k$ and exactly $\frac{(k+1)(k-2)}{2}$ chords.
To motivate further our study, let us mention that chords in cycles attracted some attention, especially in the context of bounding the chromatic number. In [MdFT13], a structure theorem for graphs where no cycle has a chord is described. A stronger theorem (where only cycles with a unique chord are excluded) is presented in [TV10], and it is generalized to graphs where no cycle of length at least 5 has a unique chord in [TP18]. All these structural descriptions imply upper bounds on the chromatic number of the graphs under consideration. In [AB15], it is shown that graphs with no cycle with exactly two chords are 6-colorable, while graphs with no cycle with exactly three chords satisfy bounds on their chromatic number. Also bounds on the chromatic number of graphs containing no cycles with exactly $k$ chords are conjectured, and they are proved in [LLP22] for sufficiently large $k$. We contribute to this line of research with the following corollary of 1. A graph is $k$-degenerate if all its subgraphs contain a vertex of degree at most $k$.
Corollary 2.
For all integers $\ell\geq 0$, if every cycle of a graph $G$ has less than $\ell$ chords, then $G$ is $\left\lceil\frac{-1+\sqrt{9+8\ell}}{2}\right\rceil$-degenerate and therefore $\left\lceil\frac{1+\sqrt{9+8\ell}}{2}\right\rceil$-colorable.
Proof.
Suppose for a contradiction that $G$ is not $\left\lceil\frac{-1+\sqrt{9+8\ell}}{2}\right\rceil$-degenerate. So, $G$ has a subgraph with minimum degree at least $\lceil k\rceil$, where
$k=\frac{-1+\sqrt{9+8\ell}}{2}+1=\frac{1+\sqrt{9+8\ell}}{2}.$
Hence, by 1, $G$ contains a cycle whose number of chords is at least:
$\frac{(\lceil k\rceil+1)(\lceil k\rceil-2)}{2}\geq\frac{(k+1)(k-2)}{2}=\ell.$
This contradiction proves the claim about the degeneracy and the claim about colorability follows. ∎
Note that for $\ell=0$, 2 restates a well-known fact: every forest is 1-degenerate. For $\ell=1$, it states that graphs where all cycles are chordless are 2-degenerate, a result already present in [MdFT13]. Observe also that 2 is tight for all $\ell\geq 0$. Indeed, for $k\geq 1$, set
$I_{k}=\left\{\ell\,\middle|\,\left\lceil\frac{-1+\sqrt{9+8\ell}}{2}\right\rceil=k\right\}.$
Every $\ell$ is in some $I_{k}$. A simple computation shows that
$I_{k}=\left\{\ell\,\middle|\,\frac{(k+1)(k-2)}{2}<\ell\leq\frac{(k+2)(k-1)}{2}\right\}.$
So, for any $\ell\in I_{k}$, 2 says that a graph with all cycles having less than $\ell$ chords is $k$-degenerate, and this is best possible since every cycle in $K_{k+1}$ has at most $\frac{(k+1)(k-2)}{2}<\ell$ chords, while $K_{k+1}$ is not $(k-1)$-degenerate.
Note that the following question can be entirely solved: does sufficiently large minimum degree imply the existence of a cycle with exactly $\ell$ chords? The answer is positive for $\ell=0$, as one can consider any induced cycle. Then the answer becomes negative for all $\ell=1,\dots,34$, and becomes positive again for $\ell=35$. To explain this, note that cliques only contain cycles whose number of chords are expressible as $a(a-3)/2$ for an integer $a$, and bicliques only contain cycles whose number of chords are expressible as $b^{2}-2b$ for an integer $b$. So the only candidates for $\ell$ are values both expressible as $a(a-3)/2$ and as $b^{2}-2b$. It turns out that all these candidate values give positive answers. Indeed, it has been announced in [SS21] that a graph of sufficiently large minimum degree contains $K_{a}$, $K_{b,b}$, or an induced cycle $C$ and a vertex $v$ with at least $\ell+3$ neighbors in $C$. In this latter case, $v$ together with a subpath of $C$ forms a cycle with exactly $\ell$ chords. The smallest positive candidate $\ell$ is 35, both corresponding to the number of chords in the hamiltonian cycle of $K_{10}$ and of $K_{7,7}$.
Dense Cyclic minors
Motivated by the edge contractions in 1, we say that a graph $H$ is a cyclic minor of a graph $G$ if a graph isomorphic to $H$ can be obtained from a Hamiltonian subgraph of $G$ by contracting the edges of the Hamiltonian cycle. So 1 just states that a large minimum degree guarantees a cycle with high minimum degree as a cyclic minor. By the Marcus-Tardos theorem [MT04], this cyclic minor can be further contracted in a cyclic way to form a complete bipartite graph (with some additional edges in the partite sets). In a nutshell, large minimum degree provides a cycle which can be cyclically contracted to a very dense subgraph.
In fact, by using the notion of $k$-linked subgraph [TW05], it is not very difficult to prove that a large average degree guarantees a large complete graph as a cyclic minor, so that one may define $f(k)$ as the smallest integer $\delta$ such that every graph of minimum degree at least $\delta$ contains $K_{k}$ as a cyclic minor. We can prove that $f(4)=3$, and $6\leq f(5)\leq 8$ and more generally $f(k)=O(k^{2})$. We propose the following open question.
Question 3.
Could it be that $f(k)=O(k\sqrt{\log k})$, matching the bound for normal minors from [Kos82, dlV83]?
Lollipops
Our method to produce a dense cycle relies on lollipops, that are subgraphs consisting of a path and a cycle containing a unique common vertex (a more formal definition is given below). They were first defined and used by Thomason in [Tho78]. Since this seminal paper, the so-called lollipop method has been extensively used, mostly to prove results about Hamiltonian cycles. We might cite about fifty papers citing [Tho78], but we just mention one that has the advantage of being recent, more related to our topic and containing a short survey and nice results [Tho18]. To the best of our knowledge, so far, the lollipop method has not been used to prove the existence of dense substructures as we do.
Outline of the paper
In Section 2, we formally define optimal lollipops and prove by a sequence of lemmas that the cycle of such a lollipop satisfies all the properties given in 1 (namely, the theorem follows directly from Lemma 8 and Lemma 11). In Section 3, we prove the claims about cyclic minors. In Section 4, we propose several conclusive remarks and open questions.
Definitions and notations
We mostly use standard terminology, see [Die24]. It is convenient here to view a path in a graph $G$ as a sequence of distinct vertices $P=p_{1}\dots p_{k}$ such that for all $i\in\{1,\dots k-1\}$, $p_{i}p_{i+1}\in E(G)$. Each edge $p_{i}p_{i+1}$ is an edge of $P$ and any other edge between vertices of $P$ is a chord of $P$. We use a similar terminology for a cycle, that we view as a sequence $C=c_{1}\dots c_{k}c_{1}$ of distinct vertices, where $k\geq 3$, and such that for all $i\in\{1,\dots k\}$, $c_{i}c_{i+1}\in E(G)$, with subscript taken modulo $k$. The length of a path (or cycle) is its number of edges. When $P$ is a path and $a$ and $b$ are vertices of $P$, we denote by $aPb$ the subpath of $P$ from $a$ to $b$.
We use the notation $V(X)$ to denote the set of vertices of any kind of object $X$ that has vertices (so $V(G)$ for a graph $G$, $V(P)$ for a path $P$ and so on). We denote by $N(v)$ the set of neighbors of $v$ in a graph $G$ and we use the notation $N_{X}(v)$ for $N(v)\cap V(X)$ (again, $X$ can be a graph, a cycle and so on). We denote by $d(v)$ the degree of a vertex $v$ (that is $|N(v)|$), and use the notation $d_{X}(v)$ for $|N_{X}(v)|$.
2 Lollipops and dense cycles
A lollipop $L$ with path $P$ and cycle $C$ in a graph $G$ is a pair $(P,C)$ where $P=p_{1}\dots p_{s}$ ($s\geq 1$) is a path of $G$, $C=c_{1}\dots c_{t}c_{1}$ ($t\geq 3$) is a cycle of $G$, $p_{s}=c_{1}$ and $V(P)\cap V(C)=\{c_{1}\}$; see Fig. 1 for an illustration.
$p_{1}$$p_{s-1}$$c_{1}$$c_{t}$$p_{1}$$p_{s-1}$$c_{1}$$c_{t}$
Figure 1: Two representations of a lollipop $L$.
A lollipop $L=(P,C)$ in $G$ is optimal if:
  • $L$ is vertex-wise maximal (that is no lollipop $L^{\prime}$ of $G$ is such that $V(L)\subsetneq V(L^{\prime}))$ and
  • among all lollipops on $V(L)$, $L$ has a cycle of maximum length (that is no lollipop $L^{\prime}=(P^{\prime},C^{\prime})$ where $V(L^{\prime})=V(L)$ is such that the length of $C^{\prime}$ is greater than the length of $C$).
Lemma 4.
Let $G$ be graph with minimum degree at least 2. For every path $Q$ of $G$, there exists an (optimal) lollipop that contains all vertices of $Q$.
Proof.
Let $Q^{\prime}=a\dots b$ be a vertex-inclusion-wise maximal path containing all vertices of $Q$. Since $d_{G}(b)\geq 2$ and $Q^{\prime}$ is vertex-inclusion-wise maximal, $b$ is adjacent to at least two vertices of $Q^{\prime}$, thus forming a lollipop. Hence, if we take among all the lollipops on $V(Q^{\prime})$ one that has a cycle with maximum length, then we obtain an optimal lollipop. ∎
From here on, we assume that $G$ is a graph with minimum degree $k\geq 2$ and $L=(P,C)$ is an optimal lollipop of $G$ with notation as above.
Lemma 5.
If $G[C]$ contains a Hamiltonian path with ends $c_{1}$ and $u$, then $N_{G}(u)\subseteq V(C)$.
Proof.
Suppose for a contradiction that $Q=c_{1}\dots u$ is an Hamiltonian path of $G[C]$, $uv\in E(G)$ and $v\notin V(C)$. If $v\in V(P)$, then the lollipop $L^{\prime}=(P^{\prime},C^{\prime})$ where $P^{\prime}=p_{1}Pv$ and $C^{\prime}=vPc_{1}Quv$ is a lollipop on $V(L)$ that has a cycle longer than $C$, a contradiction to the optimality of $L$. So $v\notin V(L)$. Then $L$ is not vertex-inclusion-wise maximal since by Lemma 4 the path $P^{\prime}=p_{1}Pc_{1}Quv$ is contained in a lollipop larger than $L$, a contradiction. Hence, $N_{G}(u)\subseteq V(C)$. ∎
Let us define recursively sets $\mathcal{S}_{1},\mathcal{S}_{2},\dots$ Each $\mathcal{S}_{i}$ is a set of Hamiltonian paths of $G[C]$ starting at $c_{1}$.
  • $\mathcal{S}_{1}=\{c_{1}c_{2}\dots c_{t},c_{1}c_{t}c_{t-1}\dots c_{2}\}$.
  • For all $i\geq 1$, let us define $\mathcal{S}_{i+1}$ from $\mathcal{S}_{i}$. For all paths $Q=c_{1}\dots u\in\mathcal{S}_{i}$ and all vertices $v$ such that $uv$ is a chord of $Q$, let $w$ be the neighbor of $v$ in $vQu$. If $vw$ is an edge of $C$, then add the path $c_{1}QvuQw$ to $\mathcal{S}_{i+1}$ (see Fig. 2).
$c_{1}$$v$$w$$u$$\in\mathcal{S}_{i}$$c_{1}$$v$$w$$u$$p_{1}$$v$$w$$u$$\in\mathcal{S}_{i+1}$
Figure 2: Construction of an element of $\mathcal{S}_{i+1}$ from an element of $\mathcal{S}_{i}$.
Any path $Q$ from some $\mathcal{S}_{i}$ is active. By extension, we also call active every vertex $u\neq c_{1}$ that is an end of an active path. Note that $c_{1}$ is not active. The following lemma is not needed, but we keep it since it illustrates the notion.
Lemma 6.
For all $i\geq 1$, $\mathcal{S}_{i}\subseteq\mathcal{S}_{i+1}$.
Proof.
We make an induction on $i$. For $i=1$, due to symmetry, is suffices to show that $c_{1}c_{t}c_{t-1}\dots c_{2}\in\mathcal{S}_{2}$. By definition, $Q=c_{1}c_{2}\dots c_{t}\in\mathcal{S}_{1}$. Since $c_{1}c_{t}$ is a chord of $Q$, it is possible to take the Hamiltonian path $Q^{\prime}=c_{1}c_{t}c_{t-1}\dots c_{2}$, so $Q^{\prime}\in\mathcal{S}_{2}$. Hence, $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$.
For $i>1$, let $Q\in\mathcal{S}_{i}$, where $Q=c_{1}\ldots u$. So there is a path $Q^{\prime}=c_{1}\dots u^{\prime}\in\mathcal{S}_{i-1}$ such that $vu^{\prime}$ is a chord of $Q^{\prime}$ and $u$ is the neighbor of $v$ in $vQ^{\prime}u^{\prime}$. Moreover $vu$ is an edge of $C$. By the induction hypothesis, since $Q^{\prime}\in\mathcal{S}_{i-1}$, we also have $Q^{\prime}\in\mathcal{\mathcal{missing}}S_{i}$. By considering this path $Q^{\prime}$, the definition of $S_{i+1}$ then implies that $Q\in\mathcal{S}_{i+1}$. ∎
Lemma 7.
$C$ contains at least $k$ active vertices. Moreover, if $d_{C}(c_{1})<k$, then $C$ contains at least $k+1$ active vertices.
Proof.
By definition $Q=c_{1}c_{2}c_{3}\dots c_{t}\in\mathcal{S}_{1}$, and thus the vertex $c_{t}$ is active and $N_{G}(c_{t})\subseteq V(C)$ by Lemma 5. So, there exist integers $22\leq i_{1}<\dots<i_{k-2}\leq t-$ 2 such that $c_{i_{j}}\in N_{G}(c_{t})$ for $j=1,\dots,k-2$. Since $c_{t}c_{i_{j}}$ is a chord of $Q$ and $c_{i_{j}}c_{i_{j}+1}$ is an edge of $C$, we have $Q_{j}=c_{1}Qc_{i_{j}}c_{t}Qc_{i_{j}+1}\in\mathcal{S}_{2}$, so $c_{i_{j}+1}$ is active. Now, $c_{2},c_{i_{1}+1},\dots,c_{i_{k-2}+1},c_{t}$ are $k$ distinct active vertices, proving the first conclusion.
Now assume that $d_{C}(c_{1})<k$. So, $c_{1}$ cannot be adjacent to all vertices in $\{c_{2},c_{i_{1}+1},\dots,c_{i_{k-2}+1},c_{t}\}$. Hence, there exists an integer $21\leq j\leq k-$ 2 such that $c_{1}c_{{i_{j}}+1}\notin E(G)$, $c_{t}c_{i_{j}}\in E(G)$, $c_{{i_{j}}+1}$ is active and $R\in\mathcal{S}_{2}$ where $R=c_{1}Qc_{i_{j}}c_{t}Qc_{{i_{j}}+1}$. Since $c_{{i_{j}}+1}$ is active and $c_{1}c_{{i_{j}}+1}\notin E(G)$, by Lemma 5, $N_{G}(c_{{i_{j}}+1})\subseteq V(C)\setminus\{c_{1}\}$. Among the neighbors of $c_{{i_{j}}+1}$, some are in $Q^{\prime}=c_{2}Qc_{{i_{j}}-1}$, and some are in $Q^{\prime\prime}=c_{{i_{j}}+3}Qc_{t}$. So, there exist integers $k^{\prime}=|N_{G}(c_{{i_{j}}+1})\cap V(Q^{\prime})|$ and $k^{\prime\prime}=|N_{G}(c_{{i_{j}}+1})\cap V(Q^{\prime\prime})|$, $12\leq i^{\prime}_{1}<\dots<i^{\prime}_{k^{\prime}}\leq{i_{j}}-$ 1 and ${i_{j}}+3\leq i^{\prime\prime}_{1}<\dots<i^{\prime\prime}_{k^{\prime\prime}}\leq t$ such that $c_{i^{\prime}_{h}}\in N_{Q^{\prime}}(c_{{i_{j}}+1})$ for $h=1,\dots,k^{\prime}$ and $c_{i^{\prime\prime}_{h}}\in N_{Q^{\prime\prime}}(c_{{i_{j}}+1})$ for $h=1,\dots,k^{\prime\prime}$. Note that $k^{\prime}+k^{\prime\prime}\geq k-2$.
$p_{1}$$p_{s-1}$$c_{1}$$c_{i^{\prime}_{h}}$$c_{i^{\prime}_{h}+1}$$c_{i_{j}}$$c_{i_{j}+1}$$c_{t}$$c_{2}$
Figure 3: A situation from the proof of Lemma 7.
For all $h=1,\dots,k^{\prime}$, the path
$R^{\prime}_{h}=c_{1}Rc_{i^{\prime}_{h}}c_{i_{j}+1}Rc_{i^{\prime}_{h}+1}=c_{1}Qc_{i^{\prime}_{h}}c_{{i_{j}}+1}Qc_{t}c_{i_{j}}Qc_{i^{\prime}_{h}+1}$
is in $\mathcal{S}_{3}$ and shows that $c_{i^{\prime}_{h}+1}$ is active (see Fig. 3).
For all $h=1,\dots,k^{\prime\prime}$, the path
$R^{\prime\prime}_{h}=c_{1}Rc_{i^{\prime\prime}_{h}}c_{{i_{j}}+1}Rc_{i^{\prime\prime}_{h}-1}=c_{1}Qc_{i_{j}}c_{t}Qc_{i^{\prime\prime}_{h}}c_{{i_{j}}+1}Qc_{i^{\prime\prime}_{h}-1}$
is in $\mathcal{S}_{3}$ and shows that $c_{i^{\prime\prime}_{h}-1}$ is active (see Fig. 4).
$p_{1}$$p_{s-1}$$c_{1}$$c_{i_{j}}$$c_{i_{j}+1}$$c_{i^{\prime\prime}_{h}-1}$$c_{i^{\prime\prime}_{h}}$$c_{t}$$c_{2}$
Figure 4: Another situation from the proof of Lemma 7.
Hence, $c_{2},$ $c_{i^{\prime}_{1}+1},$ $\dots$,… , $c_{i^{\prime}_{k^{\prime}}+1},$ $c_{{i_{j}}+1},$ $c_{i^{\prime\prime}_{1}-1},$ $\dots$,… , $c_{i^{\prime\prime}_{k^{\prime\prime}}-1},$ $c_{t}$, are $k^{\prime}+k^{\prime\prime}+3\geq k+1$ distinct active vertices, proving the second conclusion. ∎
Lemma 8.
$C$ contains at least $k+1$ vertices of degree at least $k$ (in $C$). In particular, $C$ has at least $\frac{(k+1)(k-2)}{2}$ chords.
Proof.
Indeed, by Lemma 7, $C$ contains at least $k+1$ active vertices if $d_{C}(c_{1})<k$. Moreover, if $d_{C}(c_{1})\geq k$, then $C$ has $k$ active vertices. In both cases, by Lemma 5, $C$ has at least $k+1$ vertices of degree at least $k$ in $C$. The fact that $C$ has at least $\frac{(k+1)(k-2)}{2}$ chords follows since every one of the $k+1$ vertices provides $k-2$ chords of $C$ and the number is divided by two to avoid double counting. ∎
An edge of $C$ whose both ends are non-active is passive.
Lemma 9.
All active paths go through all passive edges of $C$.
Proof.
Let $Q$ be an active path and $e$ be a passive edge of $C$. By definition of active paths, there exist an integer $i\geq 1$ such that $Q\in\mathcal{S}_{i}$. Let us prove by induction on $i$ that $Q$ goes through $e$.
If $i=1$, then $Q$ is one of $c_{1}c_{2}\dots c_{t}$ or $c_{1}c_{t}c_{t-1}\dots c_{2}$. Note that $e\notin\{c_{1}c_{2},c_{1}c_{t}\}$ because $c_{2}$ and $c_{t}$ are both active. So $Q$ goes through all passive edges of $C$, in particular through $e$.
Now suppose $i>1$. Since $Q=c_{1}\ldots u\in\mathcal{S}_{i}$, by definition of $\mathcal{S}_{i}$, there exists an active path $Q^{\prime}=c_{1}\dots u^{\prime}\in\mathcal{S}_{i-1}$ such that $Q^{\prime}$ has a chord $u^{\prime}v$ and $u$ is the neighbor of $v$ in $vQ^{\prime}u^{\prime}$. By the induction hypothesis, $Q^{\prime}$ goes through $e$. So does $Q$, since $e\neq uv$ (because $u$ is active). ∎
Lemma 10.
If $R$ is a subpath of $C$ containing only passive edges and $u$ is an active vertex, then $u$ has at most one neighbor in $R$. Moreover, such a neighbor is an end of $R$.
Proof.
Since $u$ is active, there exists a path $Q=c_{1}\dots u$ in some $\mathcal{S}_{i}$. By Lemma 9, $R$ is a subpath of $Q$, so let $a_{1},a_{2},\ldots,a_{r}=V(R)$ be the consecutive vertices of $Q$, ordered in a such a way that $c_{1},a_{1},a_{2},\ldots,a_{r},u$ appear in this order along $Q$. If $u$ is adjacent to $a_{j}$, where $1\leq j<r$, then $ua_{j}$ is a chord of $Q$ and $a_{j}a_{j+1}$ is an edge of $C$. Hence, the path $Q^{\prime}=c_{1}Qa_{j}uQa_{j+1}$ is in $\mathcal{S}_{i+1}$ and the vertex $a_{j+1}$ is active, a contradiction to $a_{j}a_{j+1}$ being passive. So, $a_{r}$ is the only possible neighbor of $u$ in $R$. ∎
Lemma 11.
There exist sets $X_{1},X_{2}\subseteq E(C)$ so that for $i\in\{1,2\}$, the graph $G_{i}$ obtained by deleting all vertices not in $C$ and contracting the edges of $X_{i}$ has at least $k$ vertices and
  • minimum degree at least $\left\lceil\frac{k+2}{2}\right\rceil$ if $i=1$, and
  • average degree at least $\tfrac{2}{3}(k+1)$ if $i=2$.
Proof.
By Lemma 7, $C$ has at least $k$ active vertices. Let $X_{0}$ be the set of passive edges of $C$. Let $G_{0}$ be the graph obtained from $G$ by deleting the vertices outside of $C$ and contracting the edges of $X_{0}$; and let $C_{0}$ be the cycle in $G$ obtained from $C$ by this contraction. By Lemma 5 and Lemma 10, the active vertices have the same degree in $G_{0}$ as in $G$.
Note that the cycle $C_{0}$ is edgewise partitioned into edges whose both ends are active vertices, and paths of length 2 whose both ends are active and whose unique internal vertex is not. Let $a_{1}b_{1}c_{1}$, …, $a_{p}b_{p}c_{p}$ be the paths of length 2 of $C_{0}$ such that the $a_{i}$’s and $c_{i}$’s are active and the $b_{i}$’s are not. Suppose that $a_{1},b_{1},c_{1}$, …, $a_{p},b_{p},c_{p}$ appear in this order along $C_{0}$ (note that possibly $c_{i}=a_{i+1}$, subscript taken modulo $p$).
Let $X^{\prime}_{1}=\{a_{i}b_{i}:i\in\{1,\ldots,p\}\}$ and $X_{1}=X_{0}\cup X^{\prime}_{1}$; thus, $G_{1}$ is the graph obtained from $G_{0}$ by contracting the edges in $X^{\prime}_{1}$. Note that the edges of $X^{\prime}_{1}$ are vertex-disjoint. Let $C_{1}$ be the cycle in $G_{1}$ obtained from $C_{0}$ by this contraction. Active vertices are in one-to-one correspondence to the vertices of $C_{1}$; in particular, $C_{1}$ has length at least $k$. Moreover, each vertex of $C_{1}$ has its two neighbors along $C_{1}$, and it is incident with at least $\left\lceil\frac{k-2}{2}\right\rceil$ chords (because contracting the edges of $X^{\prime}_{1}$ decreases the number of incident chords at most by half). Therefore, $G_{1}$ has minimum degree at least $\left\lceil\frac{k+2}{2}\right\rceil$.
We let $X_{2}=X_{1}$ if $G_{1}$ has average degree greater than $G_{0}$ and $X_{2}=X_{0}$ otherwise. Thus $G_{2}=G_{1}$ if $X_{2}=X_{1}$, and $G_{2}=G_{0}$ otherwise. Let $n_{a}$ be the number of edges in $E(G_{0})\setminus E(C_{0})$ with both active ends, let $n_{b}$ be the number of edges in $E(G_{0})\setminus E(C_{0})$ with exactly one active end, and let $m$ be the number of active vertices. Since each active vertex is incident with at least $k-2$ chords in $G_{0}$, we have
$2n_{a}+n_{b}\geq(k-2)m.$
The average degree of $G_{0}$ is at least
$\frac{2|E(G_{0})|}{|V(G_{0})|}\geq\frac{2|C_{0}|+2n_{a}+2n_{b}}{|C_{0}|}=2+\frac{2n_{a}+2n_{b}}{|C_{0}|}\geq 2+\frac{n_{a}+n_{b}}{m}.$
The graph $G_{1}$ has at least $n_{a}$ chords and $m$ vertices, and thus it has average degree at least
$\frac{2|E(G_{1})|}{|V(G_{1})|}\geq 2+\frac{2n_{a}}{m}.$
Hence, $G_{2}$ has average degree at least
$2+\frac{\max(n_{a}+n_{b},2n_{a})}{m}\geq 2+\frac{\max((k-2)m-n_{a},2n_{a})}{m}\geq 2+\tfrac{2}{3}(k-2)=\tfrac{2}{3}(k+1).$
3 Cyclic minors
In Lemma 11, we showed that the lollipop cycle $C$ (obtained in the proof of Theorem 1) admits a cyclic minor $M$ with vertices $x_{1},x_{2},\dots,x_{m}$ (enumerated in the cyclic order) with average degree more than $32k/$ 3. Let us push this argument to show that we can further contract $M$ in a cyclic way to obtain a complete bipartite graph with size $f(k)$ (where $f$ grows as a logarithm). To see this, form the 0/1 adjacency matrix $A$ of the graph $M$ in which the columns and rows are ordered according to $x_{1},x_{2},\dots,x_{m}$. The $m\times m$ matrix $A$ has more than $32km/$ 3 entries 1. By the Marcus-Tardos theorem (see [MT04]), there is a partition $X_{1},\dots,X_{f(k)},Y_{1},\dots,Y_{f(k)}$ of $x_{1},x_{2},\dots,x_{m}$ into consecutive intervals such that there is an edge between every pair of sets of vertices $X_{i},Y_{j}$, where $1\leq i,j\leq f(k)$. In other words, contracting each subset $X_{i}$ and $Y_{j}$ gives a cyclic minor of $M$ which contains a complete $(f(k)\times f(k))$-bipartite graph. A natural question is then to try to find complete graphs as cyclic minors of $C$. Let us start our investigation with small cliques as cyclic minors:
Lemma 12.
Let $C$ be a Hamiltonian cycle of a graph $F$. If $F$ has minimum degree at least $3$, then we can contract edges of $C$ to obtain $K_{4}$. If $F$ has average degree at least $6$, then we can contract edges of $C$ to obtain $K_{5}$.
Proof.
Suppose first that $F$ has minimum degree at least three. Choose a chord $uv$ of $C$ and a subpath $P$ of $C$ with ends $u$ and $v$ so that the path $P$ is as short as possible. Let $z\in V(P)\setminus\{u,v\}$ be an arbitrary vertex. Since $F$ has minimum degree at least three, $z$ is incident with a chord $zx$, and by the minimality of $|E(P)|$, we have $x\not\in E(P)$. Hence, there exist pairwise vertex-disjoint subpaths $P_{1}$, …, $P_{4}$ of $C$ such that $V(C)=V(P_{1})\cup\ldots\cup V(P_{4})$, $u\in V(P_{1})$, $z\in V(P_{2})$, $v\in V(P_{3})$, and $x\in V(P_{4})$. Contracting the edges of these paths turns $F$ into $K_{4}$.
Suppose next that $F$ has average degree at least $6$, and thus $|E(F)|\geq 3|V(F)|$. Without loss of generality, we can assume that for any edge $e\in E(C)$, contracting $e$ results in a graph of average degree less than $6$. Hence,
$|E(F/e)|\leq 3|V(F/e)|-1=3|V(F)|-4\leq|E(F)|-4,$
and thus $e$ is contained in at least three triangles in $F$. Let $z_{1}$, …, $z_{t}$ with $t\geq 3$ be the common neighbors of the ends of $e$ in $F$, in order along the path $C-e$. We say that the vertices $z_{2}$, …, $z_{t-1}$ are the peaks for $e$.
Let $e=uv$ be an edge of $C$, $z$ a peak of $e$, and $P$ a path in $C-e$ from $z$ to $u$ or $v$ chosen so that $P$ is as short as possible. By symmetry, we can assume that $u$ is an end of $P$. Note that $u$ and $v$ have a common neighbor $z_{1}$ in $P-z$, since $z$ is a peak for $e$. Let $v^{\prime}$ be the neighbor of $u$ in $P$ ($v^{\prime}=z_{1}$ is possible). Let $z^{\prime}$ be a peak of the edge $uv^{\prime}$. By the minimality of $E(P)$, we have $z^{\prime}\not\in V(P)$, and since $z^{\prime}$ is a peak for $uv^{\prime}$, we have $z^{\prime}\neq v$. Hence, $z^{\prime}$ is a vertex of the path $P^{\prime}=C-(V(P)\cup\{v\})$. Observe that contracting the edges of the paths $P-\{u,z\}$ and $P^{\prime}$ turns $F$ into $K_{5}$ (see Fig. 5). ∎
[…]
Figure 5: Model of a $K_{5}$ cyclic minor.
By 1, a graph of minimum degree at least three has a cyclic minor of minimum degree at least three, and a graph of minimum degree at least eight has a cyclic minor of average degree at least six. Thus, Lemma 12 has the following consequence.
Corollary 13.
$f(4)=3$ and $f(5)\leq 8$.
Note that $f(5)\geq 6$, since the icosahedron has minimum degree 5 and does not contain $K_{5}$ as a minor, much less a cyclic minor. In general, we have the following bound.
Lemma 14.
$f(k)=O(k^{2})$.
Proof.
By [TW05], there exists a constant $c>0$ such that for every integer $t$, every graph of minimum degree at least $ct$ contains a $t$-linked subgraph $F$. Note that this implies that for any distinct vertices $v_{1}$, …, $v_{t}$ of $F$, there exists a cycle in $F$ passing through $v_{1}$, …, $v_{t}$ in order: For each $i$, choose a vertex $u_{i}$ adjacent to $v_{i+1}$ (where $v_{t+1}=v_{1}$) so that the vertices $u_{1}$, …, $u_{t}$, $v_{1}$, …, $v_{t}$ are pairwise distinct. The $t$-linkedness of $F$ implies that $F$ contains pairwise vertex-disjoint paths $P_{1}$, …, $P_{t}$, where $P_{i}$ has ends $u_{i}$ and $v_{i}$ for each $i$. The concatenation of these paths gives the desired cycle.
If $G$ is a graph of minimum degree at least $ck^{2}$, consider a $k^{2}$-linked subgraph $H$ of $G$ and choose a matching $M$ of size $\binom{k}{2}$ in $H$. Label the vertices incident with $M$ by labels $v_{1}$, …, $v_{k(k-1)}$ so that letting $V_{i}=\{v_{(k-1)(i-1)+1},\ldots,v_{(k-1)(i-1)+(k-1)}\}$, for all distinct $i,j\in\{1,\ldots,k\}$, an edge of $M$ has one end in $V_{i}$ and the other end of $V_{j}$. Let $C$ be a cycle in $H$ passing through $v_{1}$, …, $v_{k(k-1)}$ in order. Contracting the edges of pairwise vertex-disjoint subpaths of $C$ containing $V_{1}$, …, $V_{k}$ gives $K_{k}$ as a cyclic minor of $G$. ∎
4 Concluding remarks and open problems
What we do here is in fact algorithmic, in the sense that we implicitly describe a polynomial-time algorithm whose input is a graph with minimum degree at least $k$ and whose output are cycles satisfying the conclusion of 1. However, finding an optimal lollipop is NP-hard, since solving it in polytime would imply finding a Hamiltonian cycle in polytime. Also, there might be exponentially many active paths in a lollipop. So we need to briefly explain our algorithmic claim.
In fact we neither need an optimal lollipop nor the set of all active paths to obtain the properties we want. What we need is a lollipop and a set of at least $k$ (sometimes $k+1$) active paths, all with distinct ends. Formally, we may compute them as follows (note that we maintain a set $\mathcal{A}$ of active vertices discovered so far):
  • $\mathcal{S}_{1}=\{c_{1}c_{2}\dots c_{t},c_{1}c_{t}c_{t-1}\dots c_{2}\}$, $\mathcal{A}=\{c_{2},c_{t}\}$.
  • For all $i\geq 1$, let us define $\mathcal{S}_{i+1}$ from $\mathcal{S}_{i}$. For each $Q=c_{1}\dots u\in\mathcal{S}_{i}$ and all vertices $v$ such that $uv$ is a chord of $Q$, let $w$ be the neighbor of $v$ in $vQu$. If $vw$ is an edge of $C$ and $w\notin\mathcal{A}$, then add the path $c_{1}QvuQw$ to $\mathcal{S}_{i+1}$ and $w$ to $\mathcal{A}$.
The proofs of the lemmas from Section 2 can now be seen as the description of a procedure $P$ whose input is a graph $G$ with minimum degree at least $k$, and a lollipop of $G$ that either outputs a better lollipop (with more vertices or a longer cycle), or the cycle described in 1. Let us explain some key steps justifying this claim. Each time the end of an active path has neighbors outside of the cycle of the lollipop, a better lollipop can be computed as explained in the proof of Lemma 5. Also, for each new active vertex, we test whether its neighborhood is actually in the cycle. All this guarantees that the computation is performed in polynomial time. Note that the proofs of Lemma 9 and Lemma 10 remain correct with these new settings.
Now, the global algorithm starts with a call to DFS (complexity of $O(n+m)$) to find the first lollipop, and then to the procedure $P$. While the call to $P$ fails to produce the cycles we want, it must be that a better lollipop is discovered, in which case we call $P$ again with the new lollipop. Since each new lollipop has either a longer cycle or more vertices, there are at most $n^{2}$ calls to $P$.
Back to the original definition of active paths, we have several questions. First, is every Hamiltonian path starting in $c_{1}$ active? The answer is negative, since in the complete graph $K_{6}$ with the Hamiltonian cycle $(c_{1},\dots,c_{6})$, the path $Q=c_{1}c_{2}c_{5}c_{4}c_{3}c_{6}$ is not active (there are 6 non active paths among the 120 paths starting by $c_{1}$). To see this, note that $Q$ could be either generated from $Q_{1}=c_{1}c_{6}c_{3}c_{4}c_{5}c_{2}$ or from $Q_{2}=c_{1}c_{2}c_{5}c_{6}c_{3}c_{4}$. But $Q_{2}$ can only be generated from $Q$, and the only possible generator for $Q_{1}$ (apart from $Q$) is $c_{1}c_{6}c_{3}c_{2}c_{5}c_{4}$ which can only be generated from $Q_{1}$. This fact leads to the two following questions:
Question 15.
Is there a simple criterion for a sequence of vertices to be an active path of $K_{n}$?
Question 16.
Is there a simple way to compute all the active vertices in the cycle of an optimal lollipop?
High minimum degree indeed provides cycles with many chords but one can wonder whether a similar statement could still hold for graphs with minimum degree 3. Since the disjoint union of many copies of $K_{4}$ does not contain a cycle with more than two chords, a natural question is to consider graphs with high girth $g$. It turns out that in the lollipop argument, the number of distinct active vertices (and thus chords) increases with respect to $g$. It is then reasonable to link the number of chords to the length of the host cycle. This leads to the following question:
Question 17.
Is there a constant $c>0$ such that every graph with minimum degree 3 contains a cycle of length $\ell$ with at least $c\ell$ chords?
We are very far from answering the previous problem, as even the following relaxation seems unclear:
Question 18.
Is there a function $f$ that tends to +$\infty$+ ∞ such that every graph with minimum degree 3 contains a cycle of length $\ell$ with at least $f(\ell)$ chords?
Acknowledgement
The authors thank Matthias Thomassé for the non active path of $K_{6}$.